A car leaves at 9 A.M. from a city and drives 45 miles per hour. At 11 A.M. a truck leaves from the city and drives at a rate of 60 miles per hour. At what time will the truck catch up to the car?

Answer by checkley79(3341)

D=RT
D=45T
D=60(T-2)
45T=60(T-2)
45T=60T-120
45T-60T=-120
-15T=120
T=-120/-15
T=8 HOURS AFTER THE CAR LEAVES THEY WILL MEET.
9 AM+8 HOURS=(17-12)=5 PM. ANSWER WHEN THEY MEET.